Let f(n) = (n + a)b and g(n) = nb, for any real constants a and b, where b > 0.

Using Θ, what is your conclusion for the relationship between f(n) and g(n)?
Answer #1
f(n) = Θ (g(n)) means there are positive constants c1, c2, and k, such that 0 ≤ c1g(n) ≤ f(n) ≤ c2g(n) for all n ≥ n0. The values of c1, c2, and n0 must be fixed for the function f and must not depend on n.

(n + a)b = Θ(nb)
=>  c1(nb) <= (n + a)b <= c2(nb)
=>  c1(nb) <= nb + ab <= c2(nb)
Let's assume c1 = 0.5 and c2 = 2
=>  c1(nb) <= nb + ab <= c2(nb)
=>  0.5(nb) <= nb + ab <= 2(nb)
=>  0.5(nb) <= nb + ab and nb + ab <= 2(nb)
=>  -ab <= 0.5 nb and ab <= nb
=>  -2ab <= nb and ab <= nb
=>  -2a <= n and a <= n
so, (n + a)b = Θ(nb) given c1 = 1/2, c2 = 2 and n0 = a